OpenStudy (anonymous):

After the fraction x plus 2 all over 3 minus the fraction x plus 1 all over 2x have been combined using the least common denominator of 6x, what is the numerator?

OpenStudy (anonymous):

(x-2)/3 - (x+1)/(2x) = [2x(x-2) - 3(x+1)]/(6x) = [2x^2-4x-3x-3]/(6x) = (2x^2-7x-3)/(6x)

OpenStudy (anonymous):

Thank you so much !!!

OpenStudy (anonymous):

How would it equal one of these? 2x2 + x - 3 2x2 + x - 1 - 2x2 + x + 6 -x2 + 2x + 6

OpenStudy (anonymous):

@Coolsector

OpenStudy (anonymous):

sorry i was wrong, (x+2)/3 - (x+1)/(2x) = [2x(x+2) - 3(x+1)]/(6x) = [2x^2+4x-3x-3]/(6x) = (2x^2+x-3)/(6x)

OpenStudy (anonymous):

it was x+2 and not x-2

OpenStudy (anonymous):

so now you can see :)

OpenStudy (anonymous):

Thank you!!! Can you help me with how to graph this? What is the graph of the function f(x) = the quantity of negative x squared minus 2 x minus 2, all over x minus 2?